OpenStudy (anonymous):

Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 Equation 2 |5x + 6| = 41 |2x + 13| = 28 Equation 1 has more solutions than equation 2. Equation 1 and Equation 2 have the same number of solutions. Equation 2 has more solutions than Equation 1. The number of solutions cannot be determined.

OpenStudy (anonymous):

can someone helpme?

OpenStudy (anonymous):

This is known as an inequality. To solve the inequality you would bring the numbers out of the "ABSOLUTE VALUES". For each inequality you will have a negative and a positive solution. Ex: |5x+6|=41 .... that would result to being 5x+6=41 and 5x+6=-41. Then you will just solve for x on each side. You will get a solution set for both. Equation 1 and 2 have the same solution set. Each inequality gives you two solution sets .... a negative and a positive.

OpenStudy (anonymous):

whats the awnser tho

OpenStudy (anonymous):

Read my answer. Its stated in the answer.